# Approaches to critical point theory via sequential and parametrized topological complexity

### v2

@@ Abstract.
The Lusternik-Schnirelmann category of a space was introduced to obtain a lower bound on the number of critical points of a $C^{1}$-function on a given manifold. Related to Lusternik-Schnirelmann category and motivated by topological robotics, the topological complexity (TC) of a space is a numerical homotopy invariant whose topological properties are an active field of research. The notions of sequential and parametrized topological complexity extend the ideas of topological complexity. While the definition of TC is closely related to Lusternik-Schnirelmann category, the connections of sequential and parametrized TC to critical point theory have not been fully explored yet. In this article we apply methods from Lusternik-Schnirelmann theory to establish various lower bounds on numbers of critical points of functions in terms of sequential and parametrized TCs. We carry out several consequences and applications of these bounds, among them a computation of the parametrized TC of the unit tangent bundles of $(4m-1)$-spheres.
@@ Introduction
Relating topological invariants of spaces to numbers of critical points of functions is one of the central ideas in topology and geometry. In the 1920s, Lusternik and Schnirelmann introduced their notion of category of a space to find a lower bound on these numbers for arbitrary $C^{1}$-functions on a closed smooth manifold. Since then, there have been many extensions and applications of Lusternik-Schnirelmann category to critical point theory, e.g. to infinite-dimensional Banach manifolds by R. Palais in [Pal66b] and to equivariant settings by T. Bartsch in [Bar93] as well as M. Clapp and D. Puppe in [CP91].
Another common approach to relate topological properties of manifolds to critical points of functions is Morse theory, see [Mil63]. While Morse theory often produces stronger bounds on the numbers of critical points, its downside is that the function under investigation needs to be of class $C^{2}$ and its critical points need to fulfill a non-degeneracy condition, which is only satisfied generically.
In recent years, invariants similar to Lusternik-Schnirelmann category have been studied in topological robotics. Motivated by the motion planning problem from robotics, M. Farber introduced the notion of topological complexity of a topological space $X$ in [Far03] and developed it further in several subsequent articles. For a path-connected topological space $X$, let $PX=C^{0}([0,1],X)$ be the path space endowed with the compact-open topology and let
@@
The definitions of topological complexity and its variants are strongly related to Lusternik-Schnirelmann category, in fact all of these concepts are special cases of the notion of sectional category of a fibration, which was introduced by A. Schwarz in [Sch66] under the name of genus. However, in contrast to Lusternik-Schnirelmann category, the relations between topological complexity and its variants on the one hand and critical point theory on the other hand has only been studied in a few instances. In [Far08, Definition 4.30], M. Farber defined the notion of a navigation function on a closed manifold $M$ to establish upper bounds on $\operatorname{\mathsf{TC}}(M)$ in terms of critical points of functions on $M\times M$. There, a function $F\colon M\times M\to\mathbb{R}$ is called a navigation function if it is non-negative, if $F^{-1}(\{0\})=\Delta_{M}$, where $\Delta_{M}\subset M\times M$ is the diagonal, and if it is a Morse-Bott function. By [Far08, Theorem 4.32], it holds for such a navigation function that
@@
In this article we generalize the aforementioned approach to topological complexity via critical point theory in several ways.
  • We follow the topological approach to Lusternik-Schnirelmann theory of Y. Rudyak and F. Schlenk from [RS03] to establish results on numbers of fixed points of self-maps homotopic to the identity. This generalizes the situation for critical points, which corresponds to the case of a time-1 map of a negative gradient flow.
  • We study upper bounds on sequential and parametrized topological complexities of sublevel sets in terms of critical points for arbitrary $C^{1}$-functions and obtain results for generalized navigation functions as special cases.
@@
Throughout the following, given a function $F\colon X\to\mathbb{R}$, we let $F^{\lambda}:=F^{-1}((-\infty,\lambda])$ denote the corresponding closed sublevel set for each $\lambda\in\mathbb{R}$. The following theorem is our most general result for sequential topological complexities. All of the terminology used in these assertions will be made precise below. For the moment, we just note that condition (D) mentioned in these statements is a topological generalization of the Palais-Smale condition.
@@
We further consider functions of the form
$F\colon X^{r}\to\mathbb{R},\qquad F(x_{1},\ldots,x_{r})=\sum_{i=1}^{r}f(x_{i}),$
where $f\colon X\to\mathbb{R}$ is a continuously differentiable function on the smooth manifold $X$. We will derive upper bounds for the sequential topological complexities $\operatorname{\mathsf{TC}}_{r}(f^{\lambda})$ of the closed sublevel set $f^{\lambda}=f^{-1}((-\infty,\lambda])$. Note that in contrast to the above results, here we obtain bounds for the sequential topological complexities of $f^{\lambda}$ instead of their subspace versions with respect to the ambient space $X$. An analogous, but technically more intricate, result is derived for sequential parametrized topological complexities. We will then apply the latter to a framework of equivariant fiber bundles in which the assumptions become more tangible.
@@
Let $m\in\mathbb{N}$ and let $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ be the unit tangent bundle of the standard $(4m-1)$-sphere $\mathbb{S}^{4m-1}\subset\mathbb{R}^{4m}$. Then for all $r\in\mathbb{N}$ with $r\geq 2$ it holds that
@@
This article is organized as follows. In section 1, we recall the topological approach to Lusternik-Schnirelmann theory from [RS03], in particular the notion of an index function, and its connection to the Palais-Smale condition. We introduce sequential subspace complexities in section 2, establish their index function properties and carry out the general consequences of the Lusternik-Schnirelmann theorem for sequential topological complexities. The consequences for critical points of $C^{1}$-functions are carried out in that section as well. These results are then applied to navigation functions in section 3 and to product functions in section 4, yielding the abovementioned consequences and applications. In section 5 we introduce sequential parametrized subspace complexities and establish a setting in which they form index functions to which the Lusternik-Schnirelmann-type approach can be applied. Section 6 focuses on the notion of vertical proportionality of a vector field. After introducing this condition, we apply the results of section 5 to differentiable functions with vertically proportional gradients and conclude with a discussion of parametrized navigation functions. In section 7, we will study the parametrized TC of restrictions of fibrations to sublevel sets of functions on their total spaces under the assumption that these are again fibrations and by applying the main results of section 6 to another special class of functions. To obtain examples of fibrations to which this approach applies, we study equivariant fiber bundles over homogeneous spaces in section 8 and establish an upper bound for the parametrized TC of equivariant fiber bundles between closed manifolds with homogeneous codomains. Finally, we apply this result to unit tangent bundles of odd-dimensional spheres and compute their parametrized topological complexities in certain cases.
@@ 1. Topological Lusternik-Schnirelmann theory
      (Subadditivity)  For all $A,B\subset X$ it holds that
      $\nu(A\cup B)\leq\nu(A)+\nu(B).$
@@
    An index function $\nu$ on $X$ is called $\varphi$-supervariant if
    $\nu(\varphi(A))\geq\nu(A)$
    for all closed subsets $A\subset X$.
@@
    The purpose of the notion of an index function is an axiomatization of the crucial properties of relative Lusternik-Schnirelmann categories. Namely, as shown in [CLOT03, Chapter 1], given a path-connected $X$ we obtain an index function on $X$ by putting
@@
The original version of the Lusternik-Schnirelmann theorem and in most of its generalizations are re-obtained from Theorem 1.4 by letting $X$ be a differentiable manifold and choose $\varphi$ as the time-$1$ map of a negative gradient flow of a $C^{1}$-function $f\colon X\to\mathbb{R}$. Then $\mathrm{Fix\;}\varphi=\operatorname{Crit}f$, so that Theorem 1.4 can be used to estimate the number of critical points of a function in one of its sublevels from below.
@@
[…]
@@
Put $h\colon M\to\mathbb{R}$, $h(x):=\frac{1}{\rho(\|\nabla F(x)\|)}$. Then
$X\colon M\to TM,\qquad X(x):=h(x)\cdot\nabla F(x),$
is a pseudo-gradient vector field for $F$: property (i) from Definition 1.5 follows from the boundedness of $X$ that is inherent in its definition. Considering property (ii) we compute for $x\in M$ with $\|\nabla F(x)\|\leq 2$ that $h(x)\leq 1$ and thus
$\|X(x)\|=h(x)\cdot\|\nabla F(x)\|\leq\|\nabla F(x)\|\leq 2\min\{\|\nabla F(x)\|,1\}.$
For $x\in M$ with $\|\nabla F(x)\|\geq 2$, we obtain $h(x)=\frac{1}{\|\nabla F(x)\|}$ and thus
$\|X(x)\|=1\leq 2\min\{\|\nabla F(x)\|,1\}.$
Concerning property (iii), we compute for each $x\in M$ that
$DF_{x}(X(x))=h(x)\cdot DF_{x}(\nabla F(x))=h(x)\cdot\|\nabla F(x)\|^{2}.$
By definition, $h(x)\geq\min\{\frac{1}{\|\nabla^{g}F(x)\|},1\}$ for each $x\in M$, so it follows that
@@
[…]
@@
Let $(M,g)$ be a complete Riemannian manifold and let $f\in C^{1}(M)$, such that $(f,g)$ satisfies the PS condition. Then $f_{\mu}\cap\operatorname{Crit}f$ has finitely many components for each $\mu\in\mathbb{R}$.
@@
The following result is a slight variation of [RS03, Proposition 9.1]. There, the authors study pseudo-gradients of the kind constructed in Example 1.7 as well as condition (C) instead of the PS condition. Their arguments can be transferred easily to the setting considered in this article. For the convenience of the reader, we will carry out the proof in detail.
@@
Let $(M,g)$ be a complete Riemannian Banach manifold, let $F\in C^{1}(M)$ be bounded from below and assume that $(F,g)$ satisfies the PS condition. Let $X\colon M\to TM$ be a pseudo-gradient vector field of $F$ and let $\phi_{1}\colon M\to M$ be the time-1-map of the flow of $-X$. Then $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D).
@@
Let $\phi\colon M\times\mathbb{R}\to M$ be the flow of $-X$ which is well-defined by Remark 1.6.(1) and satisfies $\phi(x,1)=\phi_{1}(x)$ for each $x\in M$. We derive from Remark 1.6.(3) that $\mathrm{Fix\;}\phi_{1}=\mathrm{sing}\,X=\operatorname{Crit}F$. Moreover, we compute for each $x\in M$ that
[…]
and derive from property (iii) of Definition 1.5 that
(1.1) $F(x)\geq F(\phi_{1}(x))\qquad\forall x\in M,$
where equality holds if and only if $DF_{\phi(x,t)}(X(\phi(x,t)))=0$ for all $t\in[0,1]$. By property (iii) of a pseudo-gradient, this is satisfied if and only if $x\in\operatorname{Crit}F=\mathrm{Fix\;}\phi_{1}$. Thus, $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies $(D_{1})$.
In the following let $\nabla^{g}F$ denote the gradient of $F$ with respect to $g$. Let $A\subset M$ be non-empty, such that $F$ is bounded on $A$ and such that $\inf\{F(a)-F(\phi_{1}(a))\mid a\in A\}=0$. As in the proof of [RS03, Proposition 9.1], one obtains sequences $(x_{n})_{n\in\mathbb{N}}$ in $A$ and $(t_{n})_{n\in\mathbb{N}}$ in $[0,1]$, such that
@@
Put $y_{n}:=\phi(x_{n},t_{n})$ for each $n$. We derive from (1.2) and property (iii) of a pseudo-gradient that
(1.3) $\lim_{n\to\infty}\nabla^{g}F(y_{n})=0.$
Let $C_{3}\geq 0$ with $|F(x_{n})|\leq C_{3}$ for all $n\in\mathbb{N}$. Then $F(y_{n})=F(\phi(x_{n},t_{n}))\leq F(x_{n})\leq C_{3}$ for each $n\in\mathbb{N}$. By assumption, $F$ is bounded from below, so it follows that $(F(y_{n}))_{n\in\mathbb{N}}$ is bounded. Together with (1.3), the PS condition for $(F,g)$ implies that $(y_{n})_{n\in\mathbb{N}}$ has a subsequence $(y_{n_{k}})_{k\in\mathbb{N}}$ which converges to some $z\in\operatorname{Crit}F=\mathrm{Fix\;}\phi_{1}$. We want to show that $(x_{n_{k}})_{k\in\mathbb{N}}$ converges to $z$ as well. Let $C_{1},C_{2}\in(0,+\infty)$ be given as in Definition 1.5. For each $k\in\mathbb{N}$, $\gamma_{k}\colon[0,t_{n_{k}}]\to M$, $\gamma_{k}(t)=\phi(x_{n_{k}},t)$ is a $C^{1}$-path from $x_{n_{k}}$ to $y_{n_{k}}$, so it holds that
[…]
where we used property (ii) of a pseudo-gradient . We derive from property (iii) that
$\forall x\in M:\quad\|\nabla^{g}F(x)\|\leq C_{2}^{-1}\cdot DF_{x}(X(x))\quad\vee\quad\|\nabla^{g}F(x)\|^{2}\leq C_{2}^{-1}\cdot DF_{x}(X(x)),$
and thus $\|\nabla^{g}F(x)\|\leq C_{3}\left(DF_{x}(X(x))+(DF_{x}(X(x)))^{\frac{1}{2}}\right)$, where $C_{3}:=\max\{C_{2}^{-1},\sqrt{C_{2}^{-1}}\}$. Inserting this into the above computation shows that
[…]

### v4

@@ Abstract.
The Lusternik-Schnirelmann category of a space was introduced to obtain a lower bound of the number of critical points of a $C^{1}$-function on a given manifold. Related to Lusternik-Schnirelmann category and motivated by topological robotics, the topological complexity (TC) of a space is a numerical homotopy invariant whose topological properties are an active field of research. The notions of sequential and parametrized topological complexity extend the ideas of topological complexity. While the definition of TC is closely related to Lusternik-Schnirelmann category, the connections of sequential and parametrized TC to critical point theory have not been fully explored yet. In this article we apply methods from Lusternik-Schnirelmann theory to establish several lower bounds of numbers of critical points of functions in terms of sequential and parametrized TCs. We carry out several consequences and applications of these bounds, among them a computation of the parametrized TC of the unit tangent bundles of (4m1)(4m-1)-spheres.
@@ Introduction
Relating topological invariants of spaces to numbers of critical points of functions is one of the central ideas in topology and geometry. In the 1920s, Lusternik and Schnirelmann introduced their notion of category of a space to find a lower bound of these numbers for arbitrary $C^{1}$-functions on a closed smooth manifold. Since then, there have been many extensions and applications of Lusternik-Schnirelmann category to critical point theory, e.g. to infinite-dimensional Banach manifolds by R. Palais in [Pal66b] and to equivariant settings by T. Bartsch in [Bar93] as well as M. Clapp and D. Puppe in [CP91].
Another common approach to relate topological properties of manifolds to critical points of functions is Morse theory, see [Mil63]. While Morse theory often produces stronger bounds of the numbers of critical points, its downside is that the considered function needs to be of class $C^{2}$ and its critical points need to fulfill a non-degeneracy condition, which is only satisfied generically.
In recent years, invariants similar to Lusternik-Schnirelmann (LS) category have been studied in topological robotics. Motivated by the motion planning problem from robotics, M. Farber introduced the notion of topological complexity of a topological space $X$ in [Far03] and developed it further in several subsequent articles. For a path-connected topological space $X$, let $PX=C^{0}([0,1],X)$ be the path space endowed with the compact-open topology and let
@@
The notions of topological complexity (TC) and its variants are strongly related to LS category, in fact all of these concepts are special cases of the notion of sectional category of a fibration, which was introduced by A. Schwarz in [Sch66] under the name of genus. However, in contrast to LS category, the relations between TC and its variants on the one hand and critical point theory on the other hand have only been studied in a few instances. In [Far08, Definition 4.30], M. Farber defined the notion of a navigation function on a closed manifold $M$ to establish upper bounds of $\operatorname{\mathsf{TC}}(M)$ in terms of critical points of functions on $M\times M$. There, a function $F\colon M\times M\to\mathbb{R}$ is called a navigation function if it is non-negative, if $F^{-1}(\{0\})=\Delta_{M}$, where $\Delta_{M}\subset M\times M$ is the diagonal, and if it is a Morse-Bott function. By [Far08, Theorem 4.32], it holds for such a navigation function that
@@
In this article we generalize the aforementioned approach to TC via critical point theory in several ways.
  • We follow the topological approach to LS theory of Y. Rudyak and F. Schlenk from [RS03] to establish results on numbers of fixed points of self-maps homotopic to the identity.
  • We study upper bounds of sequential and parametrized TCs of sublevel sets in terms of critical points for arbitrary $C^{1}$-functions and obtain results for generalized navigation functions as special cases.
@@
We want to point out that the interplay between critical points of Morse-Bott functions and LS-type invariants has been another fruitful line of research in recent years. In [KM11], H. Kadzisa and M. Mimura give a refined upper bound of LS category in terms of cone-decompositions induced by differentiable functions. Applying their approach to certain Morse-Bott functions, they obtain new computations of LS categories of homogenous manifolds, e.g. for complex Stiefel manifolds. Another relevant result studies the relations between critical points of function and the homotopic distance of maps, which was introduced by E. Macías-Virgós and D. Mosquera-Lois in [MVML22]. It generalizes various numerical invariants of spaces, among them TC and LS category. In [MVMLPS22, Theorem 4.5], the two aforementioned authors and M. J. Pereira Sáez have established an upper bound of the homotopic distance between two maps on a manifold $M$, which, given a Morse-Bott function $\Phi\colon M\to\mathbb{R}$, estimates the homotopic distance by the sum of the homotopic distances of the restrictions of the maps to the critical submanifolds of $\Phi$.
Throughout the following, given a function $F\colon X\to\mathbb{R}$, we let $F^{\lambda}:=F^{-1}((-\infty,\lambda])$ denote the corresponding closed sublevel set for each $\lambda\in\mathbb{R}$. The following theorem is our most general result for sequential topological complexities. All of the terminology used in these assertions will be made precise below. For the moment, we just note that condition (D) mentioned in these statements is a topological generalization of the Palais-Smale condition and that the sequential analogues of subspace complexities occurring in the following theorem have first been studied by N. Daundkar in [Dau24].
@@
We will further apply our results on sequential parametrized topological complexity to a framework of equivariant fiber bundles in which the technical assumptions become more tangible.
@@
Let $m\in\mathbb{N}$ and let $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ be the unit tangent bundle of the standard (4m1)(4m-1)-sphere $\mathbb{S}^{4m-1}\subset\mathbb{R}^{4m}$. Then for all $r\in\mathbb{N}$ with $r\geq 2$ it holds that
@@
This article is organized as follows. In section 1, we recall the topological approach to LS theory from [RS03], in particular the notion of an index function, and its connection to the Palais-Smale condition. We introduce sequential subspace complexities in section 2, establish their index function properties and carry out the general consequences of the LS theorem for sequential topological complexities. The consequences for critical points of $C^{1}$-functions are carried out in that section as well. These results are then applied to navigation functions in section 3 and to product functions in section 4, yielding the abovementioned consequences and applications. In section 5 we introduce sequential parametrized subspace complexities and establish a setting in which they form index functions to which the LS-type approach can be applied. Section 6 focuses on the notion of vertical proportionality of a vector field. After introducing this condition, we apply the results of section 5 to differentiable functions with vertically proportional gradients and conclude with a discussion of parametrized navigation functions. In section 7, we will study the parametrized TC of restrictions of fibrations to sublevel sets of functions on their total spaces under the assumption that these are again fibrations and by applying the main results of section 6 to another special class of functions. To obtain examples of fibrations to which this approach applies, we study equivariant fiber bundles over homogeneous spaces in section 8 and establish an upper bound of the parametrized TC of equivariant fiber bundles between closed manifolds with homogeneous codomains. Finally, we apply this result to unit tangent bundles of odd-dimensional spheres and compute their parametrized topological complexities in certain cases.
@@ 1. Topological Lusternik-Schnirelmann theory
      (Subadditivity)  For all $A,B\subset X$ it holds that $\nu(A\cup B)\leq\nu(A)+\nu(B).$
@@
    An index function $\nu$ on $X$ is called $\varphi$-supervariant if $\nu(\varphi(A))\geq\nu(A)$ for all closed subsets $A\subset X$.
@@
    The notion of index function axiomatizes the crucial properties of relative Lusternik-Schnirelmann category. Namely, as shown in [CLOT03, Chapter 1], given a path-connected $X$ we obtain an index function on $X$ by putting
@@
The original version of the Lusternik-Schnirelmann theorem and most of its generalizations can be re-obtained from Theorem 1.4 by letting $X$ be a differentiable manifold and choosing $\varphi$ as the time-$1$ map of a negative gradient flow of a $C^{1}$-function $f\colon X\to\mathbb{R}$. Then $\mathrm{Fix\;}\varphi=\operatorname{Crit}f$, so that Theorem 1.4 can be used to estimate the number of critical points of a function in one of its sublevels from below.
@@
Let $M$ be a paracompact Banach manifold and let $F\in C^{1}(M)$. Consider a Riemannian metric $g$ on $M$, let $\nabla^{g}F\colon M\to TM$ denote the gradient of $F$ with respect to the metric and let $\|\cdot\|$ denote the norm induced by $g$ on any tangent space of $M$. We call a vector field $X\colon M\to TM$ a pseudo-gradient vector field or just pseudo-gradient of $F$ if it has the following properties:
@@
Put $h\colon M\to\mathbb{R}$, $h(x):=\frac{1}{\rho(\|\nabla F(x)\|)}$. Then $X\colon M\to TM$, $X(x):=h(x)\cdot\nabla F(x)$, is a pseudo-gradient vector field for $F$: property (i) from Definition 1.5 follows from the boundedness of $X$ that is inherent in its definition. Considering property (ii), it follows from our choice of $\rho$ that
$\|X(x)\|\leq 2\min\{\|\nabla F(x)\|,1\}\quad\forall x\in M.$
Concerning property (iii), we derive from $h(x)\geq\min\{\frac{1}{\|\nabla^{g}F(x)\|},1\}$ that
@@
Let $(M,g)$ be a complete Riemannian Banach manifold and let $F\in C^{1}(M)$. We say that $(F,g)$ satisfies the Palais-Smale condition or PS condition if every sequence $(x_{n})_{n\in\mathbb{N}}$ in $M$ for which $(|F(x_{n})|)_{n\in\mathbb{N}}$ is bounded and which satisfies
@@
Let $(M,g)$ be a complete Riemannian manifold and let $f\in C^{1}(M)$, such that $(f,g)$ satisfies the PS condition. Then $f_{\mu}\cap\operatorname{Crit}f$ has finitely many connected components for each $\mu\in\mathbb{R}$.
@@
The following result is a slight variation of [RS03, Proposition 9.1]. There, the authors study pseudo-gradients of the kind constructed in Example 1.7 as well as condition (C) instead of the PS condition. Their arguments can be transferred easily to the setting considered in this article. For the convenience of the reader, we will carry out some details of the proof.
@@
Let $(M,g)$ be a complete Riemannian Banach manifold, let $F\in C^{1}(M)$ be bounded from below and assume that $(F,g)$ satisfies the PS condition. Let $X\colon M\to TM$ be a pseudo-gradient vector field of $F$ and let $\phi_{1}\colon M\to M$ be the time-1-map of the flow of X-X. Then $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D).
@@
Let $\phi\colon M\times\mathbb{R}\to M$ be the flow of X-X which is well-defined by Remark 1.6.(1) and satisfies $\phi(x,1)=\phi_{1}(x)$ for each $x\in M$. By a standard computation, one derives from property (iii) of Definition 1.5 that $(\phi_{1},F)$ satisfies $(D_{1})$.
Let $A\subset M$ be non-empty, such that $F$ is bounded on $A$ and such that $\inf\{F(a)-F(\phi_{1}(a))\mid a\in A\}=0$. As in the proof of [RS03, Proposition 9.1], one obtains sequences $(x_{n})_{n\in\mathbb{N}}$ in $A$ and $(t_{n})_{n\in\mathbb{N}}$ in $[0,1]$, such that
@@
Put $y_{n}:=\phi(x_{n},t_{n})$ for each $n$. We derive from (1.1) and property (iii) of a pseudo-gradient that $\lim_{n\to\infty}\nabla^{g}F(y_{n})=0$. Let $C_{3}\geq 0$ with $|F(x_{n})|\leq C_{3}$ for all $n\in\mathbb{N}$. Then $F(y_{n})=F(\phi(x_{n},t_{n}))\leq F(x_{n})\leq C_{3}$ for each $n\in\mathbb{N}$. By assumption, $F$ is bounded from below, so it follows that $(F(y_{n}))_{n\in\mathbb{N}}$ is bounded. The PS condition for $(F,g)$ thus implies that $(y_{n})_{n\in\mathbb{N}}$ has a subsequence $(y_{n_{k}})_{k\in\mathbb{N}}$ which converges to some $z\in\operatorname{Crit}F=\mathrm{Fix\;}\phi_{1}$. We want to show that $(x_{n_{k}})_{k\in\mathbb{N}}$ converges to $z$ as well. Let $C_{1},C_{2}\in(0,+\infty)$ be given as in Definition 1.5. For each $k\in\mathbb{N}$, $\gamma_{k}\colon[0,t_{n_{k}}]\to M$, $\gamma_{k}(t)=\phi(x_{n_{k}},t)$ is a $C^{1}$-path from $x_{n_{k}}$ to $y_{n_{k}}$, so it holds that
$d_{M}(x_{n_{k}},y_{n_{k}})$ $\leq\int_{0}^{t_{n_{k}}}\|\gamma^{\prime}_{k}(t)\|\,dt=\int_{0}^{t_{n_{k}}}\|X(\gamma_{k}(t))\|\,dt\leq C_{1}\int_{0}^{t_{n_{k}}}\|\nabla^{g}F(\gamma_{k}(t))\|\,dt,$
where we used property (ii) of a pseudo-gradient. We further derive from property (iii) that $\|\nabla^{g}F(x)\|\leq C_{3}\left(DF_{x}(X(x))+(DF_{x}(X(x)))^{\frac{1}{2}}\right)$ for each $x\in M$, where $C_{3}:=\max\{C_{2}^{-1},\sqrt{C_{2}^{-1}}\}$. Inserting this into the above computation shows that
$d_{M}(x_{n_{k}},y_{n_{k}})$ $\leq C_{1}C_{3}\int_{0}^{1}DF_{\phi(x_{n_{k}},t)}(X(\phi(x_{n_{k}},t)))\,dt+C_{1}C_{3}\Big(\int_{0}^{1}DF_{\phi(x_{n_{k}},t)}(X(\phi(x_{n_{k}},t)))\,dt\Big)^{\frac{1}{2}},$